2. To help pay your college expenses, you borrow P7000 and agree to repay the loan at the end of 5


years at 8% interest, compounded quarterly.


(a) What is the maturity value of the loan?


(b) How much interest are you paying on the loan?




1
Expert's answer
2022-06-08T17:17:22-0400

(a)


"A=P(1+r\/n)^{nt}"

"=7000(1+0.08\/4)^{4(5)}=P10401.63"

(b)


"I=A-P=10401.63-7000"

"=P3401.63"